Arcadiouse (arcadiouse):

Could someone help with this? I think I know the answer and how to do these type of problems. I just want to make sure.

Arcadiouse (arcadiouse):

|4x + 12| = 16

OpenStudy (heschamsaleh):

x=1 or x=-7

Arcadiouse (arcadiouse):

Explain your reasoning?

Arcadiouse (arcadiouse):

@Mathlete @Google @AloneS

OpenStudy (heschamsaleh):

sure

OpenStudy (heschamsaleh):

when we have a modulus it removes the negative value right ?

OpenStudy (heschamsaleh):

so we use the answer one time postive and the other negative |4x + 12| = 16 first 4x+12=16 4x=16-12 4x=4 x=1

OpenStudy (heschamsaleh):

secondly 4x+12=-16 4x=-16-12 4x=-28 x=-28/4 x=-7

OpenStudy (heschamsaleh):

got it bro :D ?

Arcadiouse (arcadiouse):

but the 16 wasn't negative?

Arcadiouse (arcadiouse):

Oh okay, so everytime we have a problem like |4x + 12| = 16, we do one time positive and the other negative? @heschamsaleh

OpenStudy (heschamsaleh):

yes my dear friend

Arcadiouse (arcadiouse):

Thanks! I really appreciate it.

OpenStudy (heschamsaleh):

you are welcome at any time when you see those sticks | | they are called modulus btw they are 2 right ? we do 2 operations - and + keep going :D anything else i d be happy to help